OpenStudy (anonymous):

At x = 2, (dy)/(dx) =

OpenStudy (unklerhaukus):

what is the formula for y

OpenStudy (dumbcow):

depends on what y(x) is ??

OpenStudy (anonymous):

here was the first part Let y= 6 x^5 - 9. Find A, B, and C so that (dy)/(dx) = Ax^B +C

OpenStudy (anonymous):

A = 30 B = 4. C =0

OpenStudy (dumbcow):

use power rule dy/dx = 30x^4

OpenStudy (unklerhaukus):

\[y(x)=6x^5-9\] \[{dy(x) \over dx} =30x^4\] \[{dy(2) \over dx} =30 \times 2^4=30 \times 16=480\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

I was stuck

OpenStudy (unklerhaukus):

thinking about our answer to see if it makes sense, 480 is a big number , if that is our gradient at that point it is very steep but our equation for y is proportional to x^4 which gets large very quickly, so it makes sense that the slope is so steep